OpenStudy (anonymous):

The atomic mass of naturally occurring neon is 20.18 amu. Naturally occurring neon is composed of two isotopes: Ne-20: 19.99 amu and Ne-22: 21.99 amu What is the percent abundance of each isotope?

OpenStudy (matt101):

Since there are only two isotopes, if you let the percentage of Ne-20 be x, the percentage of Ne-22 must be 1-x. Thus you get: \[19.99x+21.99(1-x)=20.18\]\[19.99x+21.99-21.99x=20.18\]\[2x=1.81\]\[x=0.905\] Therefore the percent abundance of Ne-20 is 0.905 (90.5%) while the percent abundance of Ne-22 is 1-0.905=0.095 (9.5%).